Two number differ by 5 and 3/5 of the greater number is equal to 3/4 to the smaller. Find the number.
Answer by lwsshak3(11628)

Two number differ by 5 and 3/5 of the greater number is equal to 3/4 to the smaller. Find the number
**
let x=larger number
x-5=smaller number
3x/5=3(x-5)/4
3x/5=(3x-15)/4
LCD=20
12x=15x-75
3x=75
x=25
x-5=20
larger number=25
smaller number=20